# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v3

@@ Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [27], [4], [9], [26, (3.14)] for more information on the topic.
Our main results will be presented in detail in Section 2, but here we highlight two of the more significant ones. In particular, Corollaries 4 and 5 of this article show respectively that for
$\sigma\geq 1-\frac{1}{5.5587\log t}\qquad\text{and}\qquad t\geq 1000,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 3.632\log t,\qquad\text{and}\qquad\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 9.842(\log t)^{11/12}.$
@@
Now we give a brief history of results on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Trudgian in [27] gives an explicit version of results in [26, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [26, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
@@
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [27], we take the opportunity to point out an error in that paper. In the quantity $A$ of [27, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [27]. In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
@@
Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us get an asymptotic improvement in the case of $1/\zeta(s)$. In particular, working within zero-free regions of the forms
@@
respectively. As far as we are aware, results of these types were not previously known. In this paper, we prove an explicit version of the first bound in (2). An explicit version of the third bound in (2) is given in [18].
@@ 2 Main results
Here we address matters of notation before listing some of the main results we prove.
@@ 2.1 Properties of the Riemann zeta function
[…]
@@
Currently, the best known result is due to [20], where above $H_{0}$, the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
@@
Here we present our main results, which are corollaries of the theorems in Sections 4 and 5.
@@
where $(W,Q)=(10,49.359)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
@@
Assume $RH(H_{0})$. Then in the region
$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad 5\leq t\leq H,$
@@
where $(\sigma_{0},Q_{RH(H_{0})})=(0.8,33.249)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ are given in Table 2. Furthermore, by assuming $RH$ instead of $RH(H_{0})$, the upper bound restriction on $t$ can be removed.
@@
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
where $Y(5.5587,10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
In addition, we have for $\sigma=1$ and $t\geq 2$,
@@
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
where $Y^{\prime}(5.5587,10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
In addition, we have for $\sigma=1$ and $t\geq 2$,
@@ 3 Some preliminaries
In this section, we present the results used in obtaining our estimates. For convenience, we begin by stating a few function theoretic results, which lie at the heart of our method.
Theorem 1.
[14, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
@@
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [25, §3.71]).
@@
[13, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$. Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$. Then
${\textup{Re}}\frac{f^{\prime}}{f}(a)=-\sum_{\rho_{j}}n_{j}\left(\frac{1}{r_{j}}-\frac{r_{j}}{R^{2}}\right)\cos\theta_{j}+\frac{1}{\pi R}\int_{0}^{2\pi}(\cos\theta)\log|f(a+Re^{i\theta})|\text{d}\theta.$
@@
Let $d>0$ be some real constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma\geq 1-1/W_{0}\log t$. Let $R^{\prime}>0$ be a parameter, chosen such that
$\frac{d}{\log t}\leq\frac{1}{W_{0}\log(t+R^{\prime})}.$ (3)
Then for $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq 1-\frac{d}{\log t}.$ (4)
Proof.
For $\sigma_{0}$ in the region stated in (4), we take $R=R^{\prime}$, $f(s)=\zeta(s)$, $a=s_{0}=\sigma_{0}+it_{0}$, and $\rho_{k}=s_{0}+r_{k}\exp(i\theta_{k})$ to be a zero of $\zeta(s)$ in Lemma 1 (where we will make the substitution $s_{0}\to s$ at the end of the proof). This gives us
${\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=-\sum_{\rho_{k}}n_{k}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}+\frac{1}{\pi R^{\prime}}\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta.$
Note that if a zero of $\zeta(s)$ lies on the boundary $|s-s_{0}|=R^{\prime}$, then since the zeros are separable, we can always find an arbitrarily smaller $R^{\prime}$ where the boundary is free of zeros.
For convenience, we define
$J:=\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta,$
for which we seek a lower bound.
For $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$, i.e. when $0\leq\cos\theta\leq 1$, we have
$\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\geq-\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
so in the ranges $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$ combined,
$J\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
For the remaining range $\pi/2\leq\theta\leq 3\pi/2$, noting that $-1\leq\cos\theta\leq 0$, we have
$\int_{\pi/2}^{3\pi/2}(\cos\theta)$ $\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
and hence for the whole range $\theta\in[0,2\pi]$,
$J\geq-4\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
Now Lemma 1 yields
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}-\frac{1}{\pi R^{\prime}}J,$ (5)
and since $r_{k}\cos\theta_{k}=-{\textup{Re}}(s_{0}-\rho_{k})$ and
$\frac{\cos\theta_{k}}{r_{k}}=\frac{r_{k}\cos\theta_{k}}{r_{k}^{2}}=\frac{-{\textup{Re}}(s_{0}-\rho_{k})}{|s_{0}-\rho_{k}|^{2}}=-{\textup{Re}}\left(\frac{1}{s_{0}-\rho_{k}}\right),$
the sum over all zeros in $|s_{0}-\rho|\leq R^{\prime}$ from (5) is
$\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}$ $=\sum_{|s_{0}-\rho|\leq R^{\prime}}-{\textup{Re}}\left(\frac{1}{s_{0}-\rho}-\frac{s_{0}-\rho}{R^{\prime 2}}\right)$
$=-\sum_{|s_{0}-\rho|\leq R^{\prime}}(\sigma_{0}-\beta)\left(\frac{1}{|s_{0}-\rho|^{2}}-\frac{1}{R^{\prime 2}}\right),$ (6)
where the second factor in (3) is non-negative.
Notice that if
$1-\frac{1}{W_{0}\log(t_{0}+R^{\prime})}\leq\sigma_{0},$
then by our non-vanishing condition on $\zeta(s)$, we always have that $\sigma_{0}-\beta\geq 0$ in the disc $|s_{0}-\rho|\leq R^{\prime}$. This condition is fulfilled by our choice of $R^{\prime}$ in (3). Thus the entire right-hand side of (3) is negative, and so we can discard any contribution from this.
Combining this with the bound on $J$ and (5) proves the lemma. ∎
@@
Let $\epsilon>0$ be a constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma>1/2$ and $t\leq t+\epsilon$. Then for a fixed $R^{\prime}<\epsilon$ and $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq\frac{1}{2}+\epsilon.$ (7)
@@
The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [28, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
@@
[29, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
@@
for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$ and
$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|\quad\text{for}\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|\quad\text{for }\textup{Re}(s)=b.\end{cases}$
@@
$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
@@
Let there be a real number $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}\geq 0$ be a number for which
@@
$|h(1+\delta_{r}+it)|$ $\leq\zeta(1+\delta_{r})|Q_{0}+1+\delta_{r}+it|,$
@@
$|\zeta$ $(s)|\leq\left(\zeta(1+\delta_{r})(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}$ (9)
×$\bm{\times}\left(k_{1}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}+1}t^{k_{2}}\log^{k_{3}}t\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
@@
$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t},$
and the right-hand side of (9) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
$t^{k_{2}}\log^{k_{3}}t\geq\frac{k_{1}(1+a_{0})^{k_{3}}(1+a_{1})^{k_{2}}}{\zeta(1+\delta_{r})}.$ (10)
@@
The proof is analogous to that of [28, Lemma 2.1], except that we are more precise in estimating (9) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
Note that different to the proof of [28, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (10)) in $\sigma$. Additionally, we found that further splitting the interval complicates (10) and raises the threshold of $t$ for the lemma to be valid. ∎

### v4

# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta function

@@ Abstract
In this article, we give explicit bounds of order $\log t$ for $\sigma$ close to $1$, for two quantities: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. We correct an error in the literature, and especially in the case of $|1/\zeta(\sigma+it)|$, also provide improvements in the constants. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$. The same method applied to the Korobov–Vinogradov zero-free region gives a new record: the unconditional bound $1/\zeta(\sigma+it)\ll(\log t)^{2/3}(\log\log t)^{1/4}$.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [29], [4], [9], [26, (3.14)] for more information on this.
The chief purpose of this article is to prove explicit upper bounds for $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Along the way, we also provide in §3 a collection of other improved bounds on quantities involving $\zeta(s)$, which might be useful for anyone seeking to do explicit computations in the area.
@@
Our main results will be presented in detail in §2, but let us first give a brief history of results on the two quantities $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Trudgian in [29] gives an explicit version of all the lemmas in [26, §3.9], which were originally due to Landau. This then allowed Trudgian to obtain explicit bounds on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$ of order $O(\log t)$. This was then extended in [15] and [7] to explicit bounds of order $O(\log t/\log\log t)$. However, estimates for the two quantities mentioned above tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
@@
While the bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem worse compared to [29], we take the opportunity to point out an error in that paper. In the quantity $A$ of [29, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,538.16)$ and $(12,59.49)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. This also impacts the subsequent bounds on $|1/\zeta(s)|$ ($R_{2}$ in [29]), which are derived from $R_{1}$. In this article, we have corrected the erroneous bounds and even slightly improved upon the (corrected) bounds.
Furthermore, by adapting a method of [7, §5] and being more careful in our derivation on bounds of $|1/\zeta(s)|$, we obtain improvements in some (erroneously better than they should be) bounds in [29]. In addition, when $\sigma=1$, the final assertion of Corollary 4 is significantly better than [3, Proposition A.2].
@@
Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us to get an asymptotic improvement in the case of $1/\zeta(s)$. In particular, working within zero-free regions of the forms
@@
respectively. As far as we are aware, results of these types were not previously known. In this paper, we prove an explicit version of the first bound in (2). The third bound in (2) is the best unconditional bound currently known, and an explicit version of it is given in [19].
@@ 2 Main results
Here we address matters of notation before listing some of our main results. We note that the arguments used in this article are largely those of [7].
@@ 2.1 Properties of the Riemann zeta function
In what follows, let $s=\sigma+it$, with $\sigma,\,t\in\mathbb{R}$, and define $H:=H_{0}-1/2$, where $H_{0}:=3.\,000\,175\,332\,800\cdot 10^{12}$ is the Riemann height, i.e., the height to which the Riemann hypothesis has been verified (see [22]). Our arguments will consider $t\leq H$ and $t\geq H$ as separate cases.
@@
Currently, the best known result is due to [21], where the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
@@
Here we present our main results, which are corollaries of the theorems in §4 and §5.
@@
where $(W,Q)=(10,71.220)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
@@
In the region
$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad t_{0}\leq t\leq H,$
@@
where $(\sigma_{0},\,t_{0},\,Q_{RH(H_{0})})=(0.8,14,23.759)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ and $t_{0}$ are given in Table 2. Furthermore, by assuming $RH$, the upper bound restriction on $t$ can be removed.
@@
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y\log t,$
where $(W,Y)=(13,1719)$ is admissible, and other values of $Y$ with corresponding $W$ are presented in Table 4.
In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 30.812\log t,$
while for $\sigma=1$ and $t\geq 2$,
@@
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(\log t)^{11/12},$
where $(W,Y)=(13,4904)$ is admissible, and other values of $Y^{\prime}$ with corresponding $W$ are presented in Table 5.
In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 87.725(\log t)^{11/12},$
while for $\sigma=1$ and $t\geq 2$,
@@ 3 Some preliminaries
In this section, we present the results used in obtaining our estimates. The argument is largely that of [15] and [7], in which detailed proofs of the core lemmas can be found. At the heart of the method lies the Borel–Carathéodory Theorem (see [15, Theorem 4]), by which the following two lemmas are proven. We remark that Lemma 1 is used in the process of proving Lemma 2.
Theorem 1 (Borel–Carathéodory).
[15, Theorem 14] Let $s_{0}$ be a complex number. Let $R$ be a positive number, possibly depending on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $\textup{Re}\,f(s)$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
@@
If in addition $f(s_{0})=0$, then for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
$|f^{\prime}(s)|\leq\frac{2R}{(R-r)^{2}}M.$
@@
[15, Lemma 15] Let $s_{0}$ be a complex number. Let $r$ and $\alpha$ be positive numbers (possibly depending on $s_{0}$) such that $\alpha<1/2$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq r$. Suppose further there is a number $A_{1}$ independent of $s$ such that,
$\left|\frac{f(s)}{f(s_{0})}\right|\leq A_{1},\qquad|s-s_{0}|\leq r.$
Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha r$ we have
$\left|\frac{f^{\prime}(s)}{f(s)}-\sum_{\rho}\frac{1}{s-\rho}\right|\leq\frac{4\log A_{1}}{r(1-2\alpha)^{2}},$
where $\rho$ runs through the zeros of $f(s)$ in the disc $|s-s_{0}|\leq\tfrac{1}{2}r$, counted with multiplicity.
@@
[15, Lemma 16] Let $s$ and $s_{0}$ be complex numbers with real parts $\sigma$ and $\sigma_{0}$, respectively. Let $r$, $\alpha$, $\beta$, $A_{1}$ and $A_{2}$ be positive numbers, possibly depending on $s_{0}$, such that $\alpha<1/2$ and $\beta<1$. Suppose that the function $f(s)$ satisfies the conditions of Lemma 1 with $r$, $\alpha$ and $A_{1}$, and that
$\left|\frac{f^{\prime}(s_{0})}{f(s_{0})}\right|\leq A_{2}.$
Suppose, in addition, that $f(s)\neq 0$ for any $s$ in the intersection of the disc $|s-s_{0}|\leq r$ and the right half-plane $\sigma\geq\sigma_{0}-\alpha r$. Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha\beta r$,
$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{8\beta\log A_{1}}{r(1-\beta)(1-2\alpha)^{2}}+\frac{1+\beta}{1-\beta}A_{2}.$
@@
If in the disc $|s-s_{0}|\leq r$, we have that $f(s)$ is analytic, non-zero, and satisfies $\left|\tfrac{f(s)}{f(s_{0})}\right|\leq A_{1}$, then for $0<\alpha_{0}<1$,
$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},\qquad|s-s_{0}|\leq\alpha_{0}r.$
@@
We follow a similar proof to that of Lemma 1 (also [15, Lemma 15]). Define $g(s)=\log(f(s)/f(s_{0}))$, where the logarithm branch is determined by $g(s_{0})=0$. Furthermore, since $f(s)\neq 0$ and is analytic in the disc $|s-s_{0}|\leq r$, hence $g(s)$ is analytic there.
By hypothesis, we have $|f(s)/f(s_{0})|\leq A_{1}$ in $|s-s_{0}|\leq r$, so in the same region,
$\textup{Re}(g(s))=\log\left|\frac{f(s)}{f(s_{0})}\right|\leq\log A_{1}.$
To complete the proof, notice that $g^{\prime}(s)=f^{\prime}(s)/f(s)$, and apply Theorem 1 in the two discs $|s-s_{0}|\leq\alpha_{0}r<r$ to obtain
$|g^{\prime}(s)|=\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2r\log A_{1}}{(r-\alpha_{0}r)^{2}}=\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},$
for $|s-s_{0}|\leq\alpha_{0}r$. ∎
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [30, Corollary 1], which is proven using a generalisation of the Phragmén–Lindelöf Principle, due to Trudgian [27, Lemma 3].
Unfortunately, Trudgian’s proof contains an error, the consequence of which is that his result is true only if one includes the restriction $\alpha_{2}=\beta_{2}$ (according to the notation of [27, Lemma 3]). In Lemma 3 below, we present a corrected version which includes this restriction. We refer the reader to [10] for an in-depth discussion on this error and a strategy to recovering the original result [^1] The author thanks Andrew Fiori for bringing this to our attention..
@@
Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
@@
for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$, $\alpha_{2}=\beta_{2}$, and
$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}(\log|Q+s|)^{\alpha_{2}}\quad\text{for }\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}(\log|Q+s|)^{\beta_{2}}\quad\text{for }\textup{Re}(s)=b.\end{cases}$
@@
$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}|\log(Q+s)|^{\alpha_{2}}}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}|\log(Q+s)|^{\beta_{2}}}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
@@
Let $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}>1/2$ be a number for which
@@
$|h(1+\delta_{r}+it)|$ $\leq k_{4}|Q_{0}+1+\delta_{r}+it|(\log|Q_{0}+1+\delta_{r}+it|)^{k_{3}},$
@@
$|\zeta(s)|\leq$ $\left(k_{4}\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}\log^{k_{3}}t$ (4)
$\bm{\times}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\left(k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}t^{k_{2}}\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
@@
$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t}.$
In addition, the right-hand side of (4) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
$t\geq\left(\frac{k_{4}}{k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}}\right)^{1/k_{2}}.$ (5)
@@
The proof is analogous to that of [30, Lemma 2.1], through a direct application of Lemma 3 with $a=1/2$ and $b=1+\delta_{r}$, except that we are more precise in estimating (4) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}(\sigma,Q_{0},t)$ and $1+a_{0}(\sigma,Q_{0},t)$.
Note that different to the proof of [30, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$. The decreasing condition can be verified by checking when the derivative of the right-hand side of (4) is negative, and this is what gives rise to the condition in (5). Additionally, we found that further splitting the interval complicates (5) and raises the threshold of $t$ for the lemma to be valid. ∎